OpenStudy (anonymous):

suppose z varies directly with x and interversley with y when x=6 and y=2 and z=15 find z when x=4 and y=p????

OpenStudy (aum):

\(\large z = k * \Large \frac xy\) Substitute x = 6, y = 2, z = 15 and find the constant of proportionality k. \(\large 15 = k * \Large \frac 62 = 3k\) \(k = 5 \) \(\large z = 5 * \Large \frac xy\) Substitute x=4 and y=p \(\large z = 5 * \Large \frac 4p = \frac{20}{p}\)

OpenStudy (anonymous):

thank you so much!

OpenStudy (aum):

you are welcome.